OpenStudy (please_help):

How is doing operations (adding, subtracting, multiplying, and dividing) with rational expressions similar to or different from doing operations with fractions? Can understanding how to work with one kind of problem help understand how to work another type? When might you use this skill in real life?

OpenStudy (anonymous):

It's no different.

OpenStudy (please_help):

Can you elaberate that please?

OpenStudy (anonymous):

"When might you use this in real life"... depends

OpenStudy (anonymous):

\[\frac{1}{1-x}-\frac{2}{3-x}=\frac{(3-x)(2-2x)}{(1-x)(3-x)}\]just like when working with fractions \[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\]where b and d do not evaluate to zero. All you do is simply substitute a,b,c,and d for expressions.

OpenStudy (anonymous):

oops \[\frac{1}{1-x}-\frac{2}{3-x}=\frac{(3-x)-(2-2x)}{(1-x)(3-x)}\]